# MIVE: New Design and Benchmark for Multi-Instance Video Editing

### v1

@@ D.3 User Study Details
Figure 15: Our user study interface and questionnaire form. Participants are presented with an input video with a source caption, an annotated video with a target caption, and 7 randomly ordered videos edited using our MIVE and six other SOTA video editing methods. Each instance mask in the annotated video is color-coded to correspond with its instance target caption. Participants are tasked to select the video with the Best Temporal Consistency, Best Textual Faithfulness, and Least Editing Leakage.
@@
To conduct our user study, we select 50 videos from our dataset covering diverse scenarios (varying numbers of objects per clip, numbers of instances per object class, and instance sizes). We show the statistics of the videos we used in our user study in Tab. 4. We edit the videos using our MIVE framework and the other six SOTA video editing methods, namely, ControlVideo [70], FLATTEN [11], RAVE [26], TokenFlow [16], FreSCo [62], and GAV [25]. We task 31 participants to select the method with the:
@@
D.4 Video-P2P Results
Figure 16: Video-P2P results on recursive multi-instance editing. The artifacts that accumulate when Video-P2P is used repeatedly for multi-instance editing is shown in red arrow. Our MIVE prevents this error accumulation since we do not edit the frames recursively.
@@ D.4 Video-P2P Results
E Additional Analysis and Ablation Studies
Methods GTC GTF FA LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA $\uparrow$ BP $\downarrow$
DMS Only NPS 0.9591 0.2667 0.7907 0.9460 0.2072 0.5587 0.6663 54.6597
Only LPS 0.9602 0.2645 0.7690 0.9483 0.2068 0.5716 0.6688 50.8549
LPS + NPS w/o Re-Inversion 0.9615 0.2674 0.7810 0.9485 0.2080 0.5776 0.6783 52.3240
Ours, Full 0.9604 0.2750 0.8557 0.9478 0.2138 0.6419 0.7100 54.3452
IPR No Modulation [44] 0.9642 0.2642 0.7468 0.9535 0.2060 0.5225 0.6553 50.1319
Dense Diffusion [28] 0.9611 0.2760 0.9029 0.9482 0.2136 0.6215 0.6891 59.2100
Ours, Full 0.9604 0.2750 0.8557 0.9478 0.2138 0.6419 0.7100 54.3452
@@ E Additional Analysis and Ablation Studies
Table 7: The full results (Global Scores and Local Scores) of our ablation study on Disentangled Multi-instance Sampling (DMS) and Instance-centric Probability Redistribution (IPR). LPS and NPS denotes Latent Parallel Sampling and Noise Parallel Sampling, respectively. The best and second best scores are shown in red and blue, respectively.
@@
Here, we present the Global Scores, which could not be included in the ablation study of the main paper due to space limitations. In the ablation study on DMS, our full method demonstrates better global editing faithfulness (GTF and FA) in addition to improved local faithfulness (LTF and IA). It also achieves competitive temporal consistency scores (GTC and LTC), consistent with the observations reported in the main paper. For the ablation study on IPR, the Global Scores show some deviation from the Local Scores, with Dense Diffusion achieving better global editing faithfulness performance. However, as shown in Fig. 7 of the main paper, Dense Diffusion may exhibit smoothing artifacts. These smoothing artifacts reduce the Local Scores but do not impact the Global Scores. In some cases, they may even improve the Global Scores, as Global Scores evaluate each frame with the whole word tokens of the caption. If each token is visible in any position of the frame, it can lead to higher Global Scores due to CLIP’s limitations in compositional reasoning [23, 36].
E.1 DMS Ablations
Figure 17: Ablation study on DMS: (a) Ablation on alternating LPS and NPS on all sampling steps ($T=50$). Increasing the number of LPS steps while setting the number of NPS steps to 1 produces more details as shown in (1) - (4). Increasing the number of NPS steps while setting the number of LPS steps to 1 produces cloudy artifacts and less details as shown in (5). Setting the number of LPS steps to 9 is a good tradeoff between qualitative and quantitative performance. See Tab. 8-(a) for quantitative results.
Figure 18: Ablation study on DMS: (b) Ablation on last NPS after alternating sampling. Reducing the number of alternating LPS-NPS steps during the initial sampling stage while increasing the number of NPS steps in the final stage enhances the quantitative performance of MIVE. Quantitative results are provided in Tab. 8-(b). However, we set the number of final NPS steps to only 10 to avoid degradation in object edges as shown in in (3) and (4).
Figure 19: Ablation study on DMS: (c) Ablation on re-inversion step on alternating sampling. Increasing the number of re-inversion steps $L$ in the alternating LPS-NPS steps improves quantitative performance as shown in Tab. 8-(c). We also observe better details with higher values of $L$. Ultimately, we set $L=3$ for the re-inversion in the alternating LPS-NPS stage.
Figure 20: Ablation study on DMS: (d) Ablation on re-inversion step on last NPS=10 with alternating LPS=9 & NPS=1 & re-inversion L=3. Increasing the number of re-inversion steps $L$ in the last NPS steps of the sampling improves quantitative performance (see Tab. 8-(d)). Higher number of NPS steps, however, give the edited objects an animation-like appearance. Thus, we limit the number of re-inversion steps in the final NPS stage to $L=2$.
Figure 21: Ablation study on DMS: (e) Ablation on Re-inversion step using 2D vs 3D Model. Using the 3D model for re-inversion improves the quantitative performance of our framework (see Tab. 8-(e)). Further, using the 3D model renders sharper edges as shown above. We, thus, use the 3D model during the re-inversion steps.
Figure 22: Ablation study on DMS: (f) ablation on another alternative configuration. Reducing the number of LPS steps during the alternating LPS-NPS stage to 4 improves FA and IA performance (see Tab. 8-(f)). However, setting the number of LPS steps to 9 better preserves subtle details, as shown in (2).
Method Global Scores Local Scores Leakage Scores
GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ BP $\downarrow$
(a) Ablation on Alternating LPS and NPS on All Sampling Steps (50)
(1) Alternate (50): LPS = 1 + NPS = 1 0.9608 0.2691 0.8098 0.9472 0.2090 0.5883 0.6851 55.1253
(2) Alternate (50): LPS = 4 + NPS = 1 0.9610 0.2667 0.7908 0.9481 0.2080 0.5914 0.6809 52.4647
(3) Alternate (50): LPS = 9 + NPS = 1 0.9605 0.2654 0.7779 0.9483 0.2075 0.5748 0.6713 51.4300
(4) Alternate (50): LPS = 14 + NPS = 1 0.9604 0.2649 0.7706 0.9483 0.2073 0.5743 0.6697 51.1941
(5) Alternate (50): NPS = 4 + LPS = 1 0.9598 0.2690 0.8181 0.9465 0.2081 0.5804 0.6727 54.8790
(b) Ablation on Last NPS after Alternating Sampling
(1) Alternate (50): LPS = 9 + NPS = 1; NPS = 0 0.9605 0.2654 0.7779 0.9483 0.2075 0.5748 0.6713 51.4300
(2) Alternate (40): LPS = 9 + NPS = 1; NPS = 10 0.9613 0.2673 0.7824 0.9485 0.2081 0.5786 0.6792 52.3441
(3) Alternate (30): LPS = 9 + NPS = 1; NPS = 20 0.9614 0.2688 0.7890 0.9483 0.2083 0.5752 0.6739 52.9035
(4) Alternate (20): LPS = 9 + NPS = 1; NPS = 30 0.9614 0.2680 0.7990 0.9485 0.2084 0.5743 0.6734 53.4412
(c) Ablation on Re-Inversion Step only on Alternating Sampling
(1) Alternate (40): LPS = 9 + Re-Inv L = 1 + NPS = 1; NPS = 10 0.9607 0.2697 0.8012 0.9479 0.2094 0.5930 0.6867 52.8582
(2) Alternate (40): LPS = 9 + Re-Inv L = 2 + NPS = 1; NPS = 10 0.9603 0.2712 0.8162 0.9475 0.2107 0.6180 0.6970 53.2689
(3) Alternate (40): LPS = 9 + Re-Inv L = 3 + NPS = 1; NPS = 10 0.9599 0.2724 0.8305 0.9471 0.2114 0.6233 0.6989 53.7494
(d) Ablation on Re-Inversion Step of Last NPS=10 with Alternating LPS=9 & NPS=1 & Re-Inversion L=3
(1) Alternate (40): LPS = 9 + Re-Inv L = 3 + NPS = 1; NPS = 10 + Re-Inv L = 1 0.9603 0.2740 0.8413 0.9477 0.2129 0.6315 0.7068 53.9953
(2) Alternate (40): LPS = 9 + Re-Inv L = 3 + NPS = 1; NPS = 10 + Re-Inv L = 2 (Ours, Full) 0.9604 0.2750 0.8557 0.9478 0.2138 0.6419 0.7100 54.3452
(3) Alternate (40): LPS = 9 + Re-Inv L = 3 + NPS = 1; NPS = 10 + Re-Inv L = 3 0.9606 0.2751 0.8591 0.9476 0.2143 0.6401 0.7090 54.7422
(e) Ablation on Re-Inversion Step using 2D vs 3D Model
(1) Alternate (40): LPS = 9 + Re-Inv L = 3 + NPS = 1; NPS = 10 + Re-Inv (2D) L = 2 0.9638 0.2749 0.8497 0.9505 0.2138 0.6385 0.7078 54.9499
(2) Alternate (40): LPS = 9 + Re-Inv L = 3 + NPS = 1; NPS = 10 + Re-Inv (3D) L = 2 (Ours, Full) 0.9604 0.2750 0.8557 0.9478 0.2138 0.6419 0.7100 54.3452
(f) Ablation on Another Alternative Configuration
(1) Alternate (40): LPS = 4 + Re-Inv L = 3 + NPS = 1; NPS = 10 + Re-Inv (3D) L = 2 0.9597 0.2775 0.8774 0.9467 0.2151 0.6626 0.7159 55.6459
(2) Alternate (40): LPS = 9 + Re-Inv L = 3 + NPS = 1; NPS = 10 + Re-Inv (3D) L = 2 (Ours, Full) 0.9604 0.2750 0.8557 0.9478 0.2138 0.6419 0.7100 54.3452
We conduct ablation studies on the hyperparameters of our Disentangled Multi-instance Sampling (DMS) technique. The quantitative results are summarized in Tab. 8, while the qualitative comparisons are shown in Figures 17, 18, 19, 20, and 21.
First, we determine the optimal balance between alternating Latent Parallel Sampling (LPS) and Noise Parallel Sampling (NPS) steps during sampling. Setting the number of NPS steps higher than the number of LPS steps results in cloudy artifacts and reduced details (Fig. 17-(5)). We observe improved details as we increase the number of LPS steps while keeping the number of NPS steps constant at 1 (Fig. 17-(1) to Fig. 17-(4)). Based on our findings, setting the LPS steps to 9 provides a good tradeoff between qualitative and quantitative performance (Tab. 8-(a)).
We also experimented on the final sampling steps of our DMS using only NPS. Increasing the number of NPS steps from 0 to 30 improves the quantitative performance of our method (Tab. 8-(b)). However, we observed that higher NPS values lead to objects losing their edges (Fig. 18-(3) and (4)). Consequently, we set NPS to 10 for the final steps of our DMS.
We further conduct experiments by introducing re-inversion in the alternating LPS-NPS steps, specifically after the LPS steps. Increasing the number of re-inversion steps $L$ after the LPS enhances the quantitative performance (Tab. 8-(c)). Moreover, as we increase $L$, the qualitative results demonstrate sharper details, such as the object shadows shown in Fig. 19-(3). We, thus, set the number of re-inversion steps in this stage to $L=3$.
We also incorporate re-inversion during the final NPS steps of our DMS sampling. Increasing the number of re-inversion steps quantitatively improves the performance of our method (Tab. 8-(d)). However, a higher number of re-inversion steps tends to introduce oversmoothing to the edited instances, giving them an animation-like appearance (Fig. 20). Therefore, we limit the number of re-inversion steps in the final NPS stage to $L=2$.
We further determine which between the 2D or 3D model used during the re-inversion steps yields better results. As shown in Tab. 8-(e), the 3D model enhances the quantitative performance of our framework. Further, using the 3D model produces sharper edges (Fig. 21-(2)). Therefore, we opt for the 3D model during the re-inversion steps.
Finally, we experiment with another DMS configuration by setting the number of LPS steps in the alternating LPS-NPS stage to 4. This results in a slight increase in FA and IA, with negligible differences in other quantitative metrics (Tab. 8-(f)). Additionally, we observe that setting LPS steps to 9 better preserves subtle details in the edited frames, such as object shadows (Fig. 22-(2)). Therefore, we adopt this configuration in the final model.
E.2 Analysis on IPR
@@
Our Instance-centric Probability Redistribution (IPR) is proposed based on two key observations: (i) manipulating the attention probabilities of the padding tokens $A_{i,j\in\boldsymbol{P}}$ may lead to artifacts; therefore, we avoid altering them; and (ii) increasing the $S$ token’s probability decreases editing faithfulness, whereas decreasing it and reallocating those values to the $\boldsymbol{T}$ and $E$ tokens improves editing faithfulness. Figure 23 illustrates the first observation, showing that altering the padding tokens tends to cause artifacts, e.g., both the alien face and the oven exhibit severe blurring and noise. In contrast, our proposed IPR avoids modifying the padding tokens $A_{i,j\in\boldsymbol{P}}$ and produces sharper and less noisy editing results, e.g., a sharp alien face and a less noisy oven.
The second observation concerns the redistribution of attention probability values among the $S$, $\boldsymbol{T}$, and $E$ tokens. Figure 24 illustrates the results of different redistribution approaches. We observe the following: (i) Redistributing attention probability values from both the $\boldsymbol{T}$ and $E$ tokens to $S$ (second row), as well as from the $\boldsymbol{T}$ token to $S$ (third row), reduces editing faithfulness. For example, the right washing machine appears gray and does not transform into a yellow washing machine, and the person remains human-like instead of transforming into an alien. Moreover, while the structure of the left washing machine is starting to change, it still does not resemble an oven. (ii) Redistributing attention probability values from the $E$ token to $S$ (fourth row) causes the right washing machine to transform into a yellow washing machine, but neither the left washing machine nor the person transforms into an oven or an alien, respectively. (iii) Redistributing attention probability values from the $S$ token to both the $\boldsymbol{T}$ and $E$ tokens by setting $\lambda_{S}=0.2$ (fifth row) begins to achieve more faithful editing. For instance, the human’s eyes start to resemble alien eyes, the left washing machine begins to transform into an oven, and the right washing machine starts to turn into a yellow washing machine. These first three observations highlight the importance of redistributing attention probabilities from $S$ to both $\boldsymbol{T}$ and $E$ tokens, which forms the main motivation for our proposed IPR. (iv) Both increasing $\lambda_{S}$ and utilizing our IPR to compute $\lambda_{S}$ enhance editing faithfulness. Notably, our proposed IPR with dynamic $\lambda_{S}$ achieves better editing faithfulness, e.g., improved faithfulness in both the alien’s face and oven, with knobs instead of a digital interface resembling the input.
@@
GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ BP $\downarrow$
(a) Ablation on $\lambda$
$\lambda=0.2$ 0.9605 0.2745 0.8469 0.9477 0.2130 0.6299 0.7072 53.8473
$\lambda=0.4$ (Ours) 0.9604 0.2750 0.8557 0.9478 0.2138 0.6419 0.7100 54.3452
$\lambda=0.6$ 0.9605 0.2753 0.8525 0.9476 0.2139 0.6459 0.7111 54.5119
(b) Ablation on $\lambda_{r}$
$\lambda_{r}=0.0$ 0.9599 0.2757 0.8770 0.9478 0.2137 0.6439 0.7041 54.8389
$\lambda_{r}=0.2$ 0.9599 0.2752 0.8663 0.9475 0.2137 0.6396 0.7023 54.6357
$\lambda_{r}=0.4$ 0.9610 0.2759 0.8589 0.9477 0.2140 0.6386 0.7101 54.5377
$\lambda_{r}=0.6$ (Ours) 0.9604 0.2750 0.8557 0.9478 0.2138 0.6419 0.7100 54.3452
$\lambda_{r}=0.8$ 0.9600 0.2746 0.8457 0.9479 0.2130 0.6438 0.7004 54.2150
$\lambda_{r}=1.0$ 0.9605 0.2728 0.8302 0.9484 0.2121 0.6277 0.6999 53.9180
(c) Ablation on Percentage of Sampling Step that Utilizes IPR
Applying IPR on The First $60\%$ Sampling Steps 0.9608 0.2750 0.8579 0.9485 0.2134 0.6350 0.7096 54.3023
Applying IPR on The First $80\%$ Sampling Steps (Ours) 0.9604 0.2750 0.8557 0.9478 0.2138 0.6419 0.7100 54.3452
Applying IPR on The First $100\%$ Sampling Steps 0.9577 0.2740 0.8447 0.9452 0.2128 0.6363 0.7082 54.0010
@@
E.3 IPR Ablations
Figure 25: Ablation study on IPR: (a) $\lambda$. Increasing $\lambda$ enhances details and textures, but increasing it too much to $\lambda=0.6$ may cause artifacts, e.g., the wood carvings get blurry. The best trade-off between details and emergence of artifacts is $\lambda=0.4$.
Figure 26: Ablation study on IPR: (b) $\lambda_{r}$. Decreasing $\lambda_{r}$ tends to improve overall editing faithfulness but introduces more noise due to the enhanced details, e.g., the shirt becomes more noisy when $\lambda_{r}$ is reduced. The best trade-off between editing faithfulness and noise is $\lambda_{r}=0.6$.
Figure 27: Ablation study on IPR: (c) percentage of sampling step where we apply our IPR. Increasing the IPR steps percentage generally enhances editing faithfulness, but excessively increasing the percentage ($100\%$) may introduce artifacts, e.g., the appearance of noise on the glass table. The best trade-off between editing faithfulness and emergence of artifacts for percentage sampling step is $80\%$.
@@ E.3 IPR Ablations
Figure 25 illustrates the results of varying the values of the hyperparameter $\lambda$. From these results, we observe that increasing $\lambda$ enhances details and textures. For example, the metal beaker with a rusty finish displays more intricate details as $\lambda$ increases. Similarly, the robot’s head and hand exhibit improved textures, e.g., the texture of the shirt becomes more natural. However, excessively increasing $\lambda$ may introduce artifacts, e.g., the wood carving becomes blurry. This occurs because a significant increase in $\lambda$ alters the spatial distribution of the initial attention values, leading to artifacts similar to DenseDiffusion [28]. In terms of quantitative performance, both $\lambda=0.4$ and $\lambda=0.6$ yield comparable results. However, due to the potential risk of artifacts, we use $\lambda=0.4$ as our default, even though $\lambda=0.6$ performs slightly better quantitatively.
For the second hyperparameter, $\lambda_{r}$, in our IPR, the qualitative results are shown in Fig. 26. We observe that increasing $\lambda_{r}$ tends to decrease overall editing faithfulness but preserves the original object shape. For example, the metal beaker begins to revert to its original shape with glass material, and the wood carving starts to lose its carvings. Conversely, decreasing $\lambda_{r}$ tends to improve overall editing faithfulness but introduces more noise due to the enhanced details. For instance, the shirts exhibit more noise when $\lambda_{r}$ is reduced. These observations are further supported by the quantitative comparison in Tab. 9, where reducing $\lambda_{r}$ improves editing faithfulness scores (GTF, FA, LTF, and IA) but worsens Leakage Scores (CIA and BP), likely due to the emergence of artifacts. Based on these findings, we select $\lambda_{r}=0.6$ to achieve a balance between overall editing faithfulness and the minimization of artifacts.

### v2

@@ C.4 User Study Details
Figure 13: Our user study interface and questionnaire form. Participants are presented with an input video with a source caption, an annotated video with a target caption, and 8 randomly ordered videos edited using our PRIMEdit and seven other SOTA video editing methods. Each instance mask in the annotated video is color-coded to correspond with its instance target caption. Participants are tasked to select the video with the Best Temporal Consistency, Best Textual Faithfulness, and Least Editing Leakage.
@@
To conduct our user study, we select 30 videos from our dataset covering diverse scenarios (varying numbers of objects per clip, numbers of instances per object class, and instance sizes). We show the statistics of the videos we used in our user study in Tab. 4. We edit the videos using our PRIMEdit framework and the other seven SOTA video editing methods, namely, ControlVideo [72], FLATTEN [14], RAVE [30], TokenFlow [20], FreSCo [63], GAV [28], and VideoGrain [65]. We task 32 participants to select the method with the:
@@
C.5 Attention Visualization
Figure 14: Attention weight visualization.
In Fig. 14, we present the attention weights for the editing results shown in Fig. 6 of the main paper. Our IPR reduces artifacts better than DenseDiffusion [32] by smoothing attention from the start (S) to the end of the text token (E), whereas DenseDiffusion concentrates on the text token (T), making it more prone to artifacts (Fig. 6-(c) of the main paper).
C.6 Video-P2P Results
Figure 15: Video-P2P [36] results on recursive multi-instance editing. The artifacts that accumulate when Video-P2P is used repeatedly for multi-instance editing is shown in red arrow. Our PRIMEdit prevents this error accumulation since we do not edit the frames recursively.
@@ C.6 Video-P2P Results
D Additional Analysis and Ablation Studies
Methods GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
IPR No Modulation [46] 0.9680 0.2614 0.7324 0.9564 0.2018 0.4920 0.6497 0.9007 0.0600
Dense Diffusion [32] 0.9699 0.2646 0.8272 0.9530 0.2078 0.5845 0.6774 0.8977 0.0604
Ours, Full 0.9677 0.2632 0.7707 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586
DMS Only SNS 0.9641 0.2623 0.7661 0.9503 0.2047 0.5288 0.6743 0.9063 0.0500
Only PNS 0.9670 0.2638 0.8008 0.9542 0.2044 0.5311 0.6718 0.8993 0.0585
SNS + PNS (no re-inv) 0.9665 0.2624 0.7688 0.9535 0.2051 0.5323 0.6728 0.9040 0.0525
Ours, Full 0.9677 0.2632 0.7707 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586
@@ D Additional Analysis and Ablation Studies
Table 8: The full results of our ablation study on Disentangled Multi-instance Sampling (DMS) and Instance-centric Probability Redistribution (IPR). LPS and NPS denotes Latent Parallel Sampling and Noise Parallel Sampling, respectively. The best and second best scores are shown in red and blue, respectively.
We present the Global Scores of in Tab. 8, which could not be included in the ablation study of the main paper due to space limitations.
@@
In the ablation study on IPR, the Global Scores align with the Local Scores, with DenseDiffusion achieving better global editing faithfulness performance. However, as shown in Fig. 6 of the main paper, Dense Diffusion may exhibit artifacts. These artifacts reduce the local scores but do not impact the global scores. In some cases, they may even improve the global scores, as global scores evaluate each frame with the whole word tokens of the caption. If each token is visible in any position of the frame, it can lead to higher global scores due to CLIP’s limitations in compositional reasoning [27, 38].
In the ablation study on DMS, our full method achieves the highest GTC and competitive global editing faithfulness (GTF and FA), while excelling in local metrics with a slight reduction in background preservation scores. This demonstrates the effectiveness of our approach in balancing global and local coherence while maintaining strong background preservation.
D.1 Analysis on IPR
@@
Our Instance-centric Probability Redistribution (IPR) is proposed based on two key observations: (i) manipulating the attention probabilities of the padding tokens $A_{i,j\in\boldsymbol{P}}$ may lead to artifacts; therefore, we avoid altering them; and (ii) increasing the $S$ token’s probability decreases editing faithfulness, whereas decreasing it and reallocating those values to the $\boldsymbol{T}$ and $E$ tokens improves editing faithfulness. Figure 16 illustrates the first observation, showing that altering the padding tokens tends to cause artifacts, e.g., the alien face exhibiting noise and oversaturation. In contrast, our proposed IPR avoids modifying the padding tokens $A_{i,j\in\boldsymbol{P}}$, preventing oversaturation and producing cleaner editing results.
The second observation concerns the redistribution of attention probability values among the $S$, $\boldsymbol{T}$, and $E$ tokens. Figure 17 illustrates the results of different redistribution approaches. We observe the following: (i) Redistributing attention probability values from both the $\boldsymbol{T}$ and $E$ tokens to $S$ (second row), as well as from the $\boldsymbol{T}$ token to $S$ (third row), reduces editing faithfulness. For example, the right washing machine appears gray and does not transform into a yellow washing machine, and the person remains human-like instead of transforming into an alien. Moreover, while the structure of the left washing machine is starting to change, it still does not resemble an oven. (ii) Redistributing attention probability values from the $E$ token to $S$ (fourth row) causes the right washing machine to transform into a yellow washing machine, but neither the left washing machine nor the person transforms into an oven or an alien, respectively. (iii) Redistributing attention probability values from the $S$ token to both the $\boldsymbol{T}$ and $E$ tokens by setting $\lambda_{S}=0.2$ (fifth row) begins to achieve more faithful editing. For instance, the human’s eyes start to resemble alien eyes, the left washing machine begins to transform into an oven, and the right washing machine starts to turn into a yellow washing machine. These first three observations highlight the importance of redistributing attention probabilities from $S$ to both $\boldsymbol{T}$ and $E$ tokens, which forms the main motivation for our proposed IPR. (iv) Both increasing $\lambda_{S}$ and utilizing our IPR to compute $\lambda_{S}$ enhance editing faithfulness. Notably, our proposed IPR with dynamic $\lambda_{S}$ achieves better editing faithfulness, e.g., improved faithfulness in both the alien’s face, oven, and yellow washing machine.
D.2 IPR Ablations
@@
GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
(a) Ablation on Percentage of Sampling Step that Utilizes IPR
Applying IPR on The First $10\%$ Sampling Steps (Ours) 0.9660 0.2636 0.7773 0.9528 0.2060 0.5390 0.6806 0.9055 0.0503
Applying IPR on The First $20\%$ Sampling Steps 0.9657 0.2645 0.7825 0.9524 0.2062 0.5387 0.6783 0.9053 0.0502
Applying IPR on The First $30\%$ Sampling Steps 0.9656 0.2648 0.7923 0.9516 0.2062 0.5494 0.6782 0.9052 0.0501
(b) Ablation on $\lambda$
$\lambda=0.4$ 0.9658 0.2621 0.7635 0.9532 0.2048 0.5345 0.6793 0.9053 0.0506
$\lambda=0.5$ (Ours) 0.9658 0.2623 0.7668 0.9531 0.2050 0.5365 0.6823 0.9054 0.0505
$\lambda=0.6$ 0.9658 0.2627 0.7685 0.9531 0.2053 0.5334 0.6791 0.9054 0.0504
$\lambda=0.7$ 0.9659 0.2632 0.7732 0.9530 0.2057 0.5363 0.6807 0.9055 0.0503
(c) Ablation on $\lambda_{r}$
$\lambda_{r}=0.4$ 0.9659 0.2635 0.7745 0.9530 0.2053 0.5360 0.6764 0.9052 0.0513
$\lambda_{r}=0.5$ (Ours) 0.9659 0.2625 0.7692 0.9532 0.2051 0.5349 0.6816 0.9053 0.0509
$\lambda_{r}=0.6$ 0.9658 0.2623 0.7668 0.9531 0.2050 0.5365 0.6823 0.9054 0.0505
$\lambda_{r}=0.7$ 0.9660 0.2625 0.7737 0.9530 0.2053 0.5375 0.6807 0.9055 0.0502
@@
Figure 18: Ablation study on IPR: (a) percentage of sampling steps where we apply our IPR. Increasing the IPR step percentage generally enhances editing faithfulness; however, applying it excessively can lead to the loss of object structures (e.g., the alien’s face and bag strap). Our results indicate that applying IPR for 10% of the sampling steps provides the best trade-off between editing faithfulness and structural preservation.
Figure 19: Ablation study on IPR: (b) $\lambda$. Increasing $\lambda$ steers the editing more towards the edit caption, but increasing it too much to $\lambda=0.7$ may cause oversaturation, e.g., the alien’s face and clothes. The best trade-off is $\lambda=0.5$.
Figure 20: Ablation study on IPR: (c) $\lambda_{r}$. Increasing $\lambda_{r}$ enhances details (e.g., the alien’s face) but the object tends to revert back to the original object. The best trade-off between editing faithfulness and enhancement in details is $\lambda_{r}=0.5$.